# Sparse Subnetwork Enhancement for Underrepresented Languages in Large Language Models

### v1

@@ 4.1 Training Setup
We conduct experiments with Llama-3.1-8B-base and Mistral-Nemo-12B-base Grattafiori et al. (2024), using GlotCC Kargaran et al. (2024) as the source of target-language training data. By default, each language subnetwork is trained on up to 100M tokens in the target language, with additional experiments varying the training size between 50M and 200M tokens for a subset of languages. We focus on 12 underrepresented languages Joshi et al. (2020), reported in Table 1.
Subnetworks are identified for each language using 100MB of GlotCC data, with thresholds $K=5\%$ and activity thresholds of 0.95, following Gurgurov et al. (2025a). Each subnetwork is fine-tuned for one epoch with batch size 2, a learning rate of $1\times 10^{-4}$, and weight decay of 0.01. We use AdamW Loshchilov and Hutter (2019) with gradient clipping Mikolov et al. (2012) for stability. Validation is performed every 500 steps, and the best checkpoint is selected by validation loss on the corresponding Flores-200 subset Costa-Jussà et al. (2022).
Lang. # Neurons # Trainable Param-s %
@@
Table 1: Neuron, parameter counts for different language subnetworks, and percentages of model being trained within Llama-3.1-8B. The statistics for Mistral-Nemo-12B are available in Appendix A.
@@ 4.2 Evaluation Tasks
Target-language performance. We assess translation quality using the "valtest" splits of Flores-200 Costa-Jussà et al. (2022) and the "test" splits of OPUS-100 Tiedemann (2012), reporting results in Sentence BLEU Post (2018). For comprehension and classification, we evaluate on Belebele Bandarkar et al. (2024) and SIB-200 Adelani et al. (2023), using accuracy as the primary metric.
General capabilities. To measure the impact of adaptation on overall model performance, we report accuracy on several standard benchmarks before and after fine-tuning: MMLU Hendrycks et al. (2021), Hellaswag Zellers et al. (2019), PIQA Bisk et al. (2019), Winogrande Sakaguchi et al. (2019), and ARC (Easy and Challenge) Clark et al. (2018).
For all tasks, we use deterministic generation with zero temperature to eliminate variance introduced by stochastic decoding. The prompts and hyperparameters specific to each evaluation are listed in Appendix C.2.
Figure 2: Neuron distributions over the layers of Llama-3.1-8B for all 12 languages. The distributions for Mistral-Nemo-12B are in Appendix A.
Metric Base Model Full FT FFN-Only FT LoRA Random Neurons Target Neurons
Target Language Performance
FLORES tg→en 28.82 11.44 25.06 28.00 30.95 31.37
FLORES en→tg 11.62 9.29 13.70 15.60 15.84 16.57
BELEBELE 66.52 21.22 38.35 56.61 59.82 66.84
SIB200 40.89 11.23 21.65 33.37 34.07 44.94
OPUS tg→en 19.59 10.67 20.01 22.68 24.52 24.34
OPUS en→tg 13.32 7.71 13.04 16.33 16.93 17.23
Avg. 29.56 11.88 21.81 28.41 29.89 33.02
General Capabilities
ARC challenge 70.90 20.40 32.94 59.45 63.69 70.62
ARC easy 85.79 23.86 46.77 77.12 80.83 84.72
Hellaswag 40.38 13.49 22.56 35.02 34.90 40.19
PIQA 71.87 34.84 48.19 64.68 69.69 72.55
Winogrande 51.40 38.15 48.34 50.76 51.24 52.41
MMLU 58.06 22.34 32.59 49.59 52.45 57.52
Avg. 63.07 25.51 38.57 56.10 58.80 63.00
Table 2: Average results for enhancing language subnetworks within Llama-3.1-8B, evaluated on both monolingual capabilities and general performance. All models were trained on 100M tokens in the target language. The downstream results for the individual languages are in Appendix C.
@@ 4.3 Baselines
We compare our sparse fine-tuning approach against four baselines:
  • (i) full model fine-tuning,
  • (ii) fine-tuning only the FFN components,
  • (iii) fine-tuning a random subset of FFN neurons of the same size as the LAPE-selected subnetwork, and
  • (iv) fine-tuning a LoRA Hu et al. (2021) applied to gate_proj and gate_up, configured to match the average parameter count used for most languages, with rank set to 62 and alpha set to 32, resulting in 57M trainable parameters (0.7%).
For full-model and FFN-only fine-tuning, we set the learning rate to $1\times 10^{-5}$ to account for the larger number of trainable parameters, while keeping all other hyperparameters identical to those of our method. For LoRA and the random-subset baseline, we fine-tune with exactly the same hyperparameters as our method.
@@ 5 Experimental Results
In this section, we present our experimental findings. We first report the results of language-specific subnetwork identification, analyzing the extent to which distinct subnetworks emerge across languages. We then evaluate the effectiveness of subnetwork fine-tuning, highlighting its impact on performance relative to baselines.
@@ 5.1 Subnetwork Identification
Table 1 summarizes the number of language-specific neurons, the corresponding trainable parameters, and the fraction of the model trained for each language-specific subnetwork within Llama-3.1-8B. The table shows substantial variation across languages: for instance, Welsh and Nepali subnetworks are relatively large, covering over 0.7% of the model parameters, whereas Georgian and Macedonian subnetworks are much smaller, accounting for less than 0.3%. This variation might reflect differences in language complexity and the amount of language-specific signal captured by GlotCC. The subnetwork sizes for Mistral-Nemo-12B are in Appendix A.
Figures 2 and 11 depict the distribution of the subnetwork neurons across layers for all 12 languages. We observe that most language-specific neurons are concentrated in the upper layers, and related languages exhibit moderate overlap in their neuron sets (such as Slovak and Slovenian, or Latvian and Lithuanian–see Appendix A for details). These patterns align with prior findings from Gurgurov et al. (2025a). Overall, the weights associated with these subnetworks account for roughly 0.2% to 1% of the total model parameters, highlighting the efficiency of this approach.
@@ 5.2.1 Average Task Performance
Table 2 reports average downstream task performance across the 12 target languages for different fine-tuning configurations, along with results on general-purpose benchmarks for Llama-3.1-8B. Targeted sparse fine-tuning of language-specific subnetworks consistently yields the largest gains on target-language tasks, outperforming full-model, FFN-only, LoRA, and random-subset baselines, while maintaining general capabilities as measured by MMLU and commonsense tasks. For example, on FLORES, targeted subnetworks achieve average improvements of up to 5 points, while preserving or slightly improving performance on general benchmarks such as MMLU and Winogrande. This contrasts with the findings of Mondal et al. (2025), who also employed LAPE for neuron identification but restricted their analysis to only three relatively high-resource languages and adapted LoRA modules on top of FFNs rather than directly fine-tuning neurons inside the model, leading to more limited improvements in cross-lingual transfer. By contrast, full-model and FFN-only fine-tuning not only fail to improve target-language performance but also substantially degrade general capabilities. The random neuron-subset and LoRA baselines show intermediate behavior, with modest target-language gains but noticeable reductions in overall performance, underscoring the advantage of precisely selecting language-specific subnetworks for sparse adaptation. Results for individual languages for both Llama-3.1-8B and Mistral-Nemo-12B are provided in Appendix C.
@@
5.2.2 Language-Wise Variation
As shown in Figure 3, the magnitude of improvement varies considerably across languages. Languages such as Afrikaans, Maltese, and Welsh benefit the most from targeted fine-tuning, with average gains exceeding 5 points across all tasks, whereas languages such as Latvian and Macedonian see only marginal improvements. A likely explanation is that higher-resource languages are already well represented in pre-training, limiting the marginal utility of additional fine-tuning data. In some cases, the model may even have encountered overlapping material during pre-training, further reducing the benefit of adaptation Lee et al. (2021). By contrast, lower-resource languages, on which the base model underperforms, gain substantially more from precise subnetwork adaptation. The only decreases we observe are for Georgian with Llama and Slovenian and Nepali with Mistral. Since both models show improvements on these languages overall (i.e., Mistral for Georgian and Llama for Slovenian and Nepali), and improvements are also evident on individual tasks (here we report averages), the gains for these languages are clearly possible but may not materialize due to the relatively small number of subnetwork parameters allocated to them. Finally, we note that some languages–particularly highly morphological ones or those with different scripts–do improve under our methods, but less than languages such as Afrikaans and Maltese. This may indicate that they require larger or smaller subnetworks (i.e., higher or lower K%). We leave the systematic exploration of optimal subnetwork sizes for different languages to future work.
Figure 4: Validation loss curves for different fine-tuning configurations of Llama-3.1-8B on Afrikaans (all trained on 100M tokens). The plots for the other tested languages are in Appendix B.
Figure 5: Overall improvement scores against various data sizes for a subset of 5 languages when sparsely fine-tuning Llama-3.1-8B.
(a) Down Projection
(b) Up Projection
(c) Gate Projection
@@ 5.2.2 Language-Wise Variation
(a) Original Llama-3.1-8B
(b) Fine-tuned (100M tokens)
@@
Figure 7: Cosine similarity for post-FFN hidden state activations of the original and sparsely fine-tuned Llama-3.1-8B between Maltese (left, middle) and average across target languages (right). Full results in Appendix D.
5.2.3 Training Dynamics
Figure 4 illustrates the training dynamics for Afrikaans. Sparse fine-tuning of target-language subnetworks generally converges faster and reaches lower validation loss than random-subset, FFN-only, full-model, or LoRA fine-tuning. While full-model and FFN-only fine-tuning occasionally achieve lower loss on the target language (see Appendix B) due to the greater capacity of the updated networks, this is often accompanied by pronounced overfitting and a substantial drop in general capabilities.
5.2.4 Effect of Training Data Size
Figure 5 further examines the effect of varying training data size (50M, 100M, 150M, and 200M tokens) for a subset of seven languages (Macedonian, Latvian, Nepali, Slovak, Afrikaans, Maltese, and Lithuanian) using Llama-3.1-8B. We observe relatively consistent improvement trends with increasing data across most tested languages, irrespective of subnetwork size, suggesting that additional data could further enhance performance. This differs from the findings of Zhao et al. (2024), who fine-tune an ablation-identified subset of neurons using between 100 and 800 documents–likely comprising fewer tokens–and report saturation after 400 documents. However, for Lithuanian we similarly observe a consistent decrease with larger training token sizes, which may indicate that the additional data is not helpful Lee et al. (2021) or that the capacity of the identified subnetwork is insufficient to effectively utilize it. Due to computational constraints and the limited availability of data for some languages (e.g., Maltese and Afrikaans, with at most 100–150M tokens), we leave the exploration of larger-scale training to future work.
6 Effects of Sparse Fine-Tuning
In this section, we present additional analyses focusing on two aspects: the changes in model weights induced by our sparse fine-tuning approach and the evolution of representation similarities before and after fine-tuning.
@@ 6.1 Weight Change Analysis
Our analysis of the weight changes during fine-tuning reveals a notable asymmetry between the up/gate projections and the down projection. As shown in Figure 6, the distribution of weight changes for the gate_proj and up_proj remains relatively consistent across all layers. The small size of the box plots indicates that most weight changes are exceptionally small and tightly clustered around zero. This suggests that the fine-tuning process is highly precise, making only minor adjustments to most weights. In contrast, the weight changes in the down_proj show an increase in magnitude across all layers, with especially large shifts in the later layers. This suggests that individual weights undergo more substantial alterations there, aligning with the greater prominence of language-specific neurons in those parts of the model. This finding suggests a functional distinction in how different weight components of the FFN layers contribute to language fine-tuning.
This asymmetry can be explained by the distinct roles of these projections within the SwiGLU-based architecture Shazeer (2020). The up_proj and gate_proj work together to create a gated feature space, refining the model’s ability to select and activate an appropriate subset of existing features for the new language. Updating these weights primarily involves slightly re-weighting established feature-extraction capabilities rather than fundamentally altering the feature space itself. Conversely, the down_proj is responsible for an important integrative step, taking the activated features and synthesizing a new representation. Our observation that its weight changes are substantially larger in later layers indicates this is where the most significant language-specific re-wiring occurs, as these deeper layers learn to reorganize high-level concepts to adapt to the new language.
6.2 Representation Analysis
@@
Targeted neuron fine-tuning not only improves performance in the target language but also enhances cross-lingual alignment. To investigate this, we examine whether parallel sentences become closer in embedding space after adaptation. Figure 7 presents cosine similarities between post-layer hidden state activations of Flores-200 sentences in Maltese and several high-resource languages (Arabic, Italian, English, French, German, Spanish, Russian, Chinese, Japanese) across layers. We observe a consistent increase in similarity for all languages after adaptation, especially in the later layers, as the early layers already exhibit relatively high similarity.
@@
The average results across all languages in Figure 7(c) show that targeted sparse fine-tuning achieves the strongest cross-lingual alignment, with an average cosine similarity of 90.47 between parallel sentences. This represents a substantial improvement over all considered fine-tuning baselines. These findings indicate that selectively updating neurons identified as important for the target language produces more aligned representations across languages, suggesting that sparse adaptation can improve cross-lingual generalization without compromising overall model structure. Consistent with these results, Sundar et al. (2025) also report improved cross-lingual alignment when manipulating language-relevant activations.
7 Conclusion
We presented a framework for enhancing the capabilities of LLMs in underrepresented languages through targeted fine-tuning of language-specific subnetworks. By identifying neurons with strong language selectivity using LAPE and adapting only their associated weights, our approach achieves substantial gains in low- and mid-resource languages while preserving general-purpose performance and updating less than 1% of parameters. Beyond improved accuracy, we showed that targeted adaptation enhances cross-lingual alignment and yields favorable training dynamics. Our release of language-specific neuron identifications for over 100 languages provides a scalable and cost-effective resource for extending state-of-the-art models to underrepresented languages, helping narrow the gap in global language coverage.
Limitations
While our framework improves performance across diverse languages, it relies on accurate identification of language-specific neurons, which can be sensitive to dataset quality and language coverage. Our experiments primarily focus on mid- and low-resource languages, leaving open the question of how the method scales to extremely low-data settings or to languages that are highly typologically distant. Furthermore, we do not explore varying the capacity of identified language-specific subnetworks through the parameter K%. Understanding how many parameters are needed to effectively capture language-specific knowledge, and how much data is required to leverage them, remains an important direction for future work. Such analyses, conducted at the per-language or per-language-family level, have significant potential.

### v2

@@ 4.1 Training Setup
We experiment with Llama-3.1-8B, Mistral-Nemo-12B and Aya-Expanse-8B, using GlotCC (Kargaran et al., 2024) as target-language training data. Each subnetwork is trained on up to 100M tokens by default, with additional experiments varying the size from 50M to 200M tokens for a subset of languages. We focus on 12 underrepresented languages (Joshi et al., 2020), reported in Table 1.
Subnetworks are identified for each language using 100MB of the target data, with thresholds $\rho=5\%$ and activity thresholds of 0.95. Each subnetwork is trained for one epoch (batch size 2) using AdamW (Loshchilov & Hutter, 2019) with a learning rate of $1\times 10^{-4}$, weight decay 0.01, and gradient clipping (Mikolov et al., 2012). We validate every 500 steps and select the best checkpoint via validation loss on the corresponding Flores-200 subset (Costa-Jussà et al., 2022).
Lang. # Neurons # Param-s % Model
@@
Table 1: Language subnetwork statistics for Llama-3.1-8B: neuron counts, trainable parameters, and percentage of total model parameters.
@@ 4.2 Evaluation Tasks
Figure 2: Neuron distributions for Llama-3.1-8B.
Target-language performance. We assess translation quality on Flores-200 (“valtest”) and OPUS-100 (“test”) via Sentence BLEU (Post, 2018), while comprehension and classification are evaluated on Belebele (Bandarkar et al., 2024) and SIB-200 (Adelani et al., 2024) using accuracy.
General capabilities. We measure the effect of adaptation on general capabilities via accuracy on several benchmarks: MMLU (Hendrycks et al., 2021), Hellaswag (Zellers et al., 2019), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019), ARC (Clark et al., 2018).
For all tasks, we use deterministic generation with zero temperature to eliminate variance introduced by stochastic decoding. The prompts and hyperparameters specific to each evaluation are listed in Appendix E.
@@ 4.3 Baselines
We compare our sparse fine-tuning approach against five baselines:
  • (i) Full FT, full model fine-tuning ,
  • (ii) FFN FT, fine-tuning only the FFN components,
  • (iii) Rand. N-s, fine-tuning a random set of FFN neurons, uniformly sampled across layers, matching the size of the LAPE-selected subnetwork,
  • (iv) LoRA (Hu et al., 2022) applied to gate_proj, up_proj, and down_proj, configured to match our average parameter count (rank 62, alpha 32, yielding 57M trainable parameters, 0.7%), and
  • (v) IA3 (Liu et al., 2022) applied to the standard target modules (k_proj, v_proj, down_proj), which learns rescaling vectors rather than low-rank updates, providing approximately 0.5M trainable parameters (0.006%).
We include both LoRA and IA3 as parameter-efficient baselines: LoRA remains the most widely adopted PEFT method, while IA3 represents an extreme point on the parameter-efficiency spectrum, achieving competitive performance with orders of magnitude fewer parameters.
For full-model and FFN-only fine-tuning, we use a learning rate of $1\times 10^{-5}$ to avoid instability from the larger parameter count. For LoRA, IA3, and the random-subset baseline, we use identical hyperparameters to our method ($1\times 10^{-4}$). Consistent with Schulman & Lab (2025), we find that PEFT methods require learning rates approximately $10\times$ larger than full fine-tuning; using lower rates results in undertraining under our fixed budget.
@@ 5 Experimental Results
In this section, we present our experimental findings. We first analyze language-specific subnetwork identification, examining the extent to which distinct subnetworks emerge across languages. We then discuss subnetwork fine-tuning and its impact on performance relative to baselines.
Metric Base Full FT FFN FT LoRA IA3 Rand. N-s Lang. N-s
% Param-s tuned 100% 70% 0.7% 0.006% 0.2–1% 0.2–1%
Monolingual Capabilities
FLORES tg→en 28.82 11.44 25.06 28.00 29.61 30.95 31.37
FLORES en→tg 11.62 9.29 13.70 15.60 12.02 15.84 16.57
BELEBELE 66.52 21.22 38.35 56.61 64.33 59.82 66.84
SIB200 40.89 11.23 21.65 33.37 42.20 34.07 44.94
OPUS tg→en 19.59 10.67 20.01 22.68 21.63 24.52 24.34
OPUS en→tg 13.32 7.71 13.04 16.33 13.70 16.93 17.23
Avg. 29.56 11.88 21.81 28.41 30.05 29.89 33.02
General Performance
ARC challenge 70.90 20.40 32.89 59.62 70.57 63.74 70.62
ARC easy 85.79 23.86 47.02 77.00 84.78 80.81 84.74
Hellaswag 40.38 13.48 22.53 34.99 39.80 34.90 40.21
PIQA 71.87 34.85 48.21 64.64 74.15 69.70 72.61
Winogrande 51.47 38.06 48.45 50.88 52.07 51.27 52.35
MMLU 58.07 22.32 32.42 49.47 57.83 52.49 57.52
Avg. 63.08 25.50 38.59 56.10 63.20 58.82 63.01
Table 2: Average results across 12 languages for Llama-3.1-8B. green: best; grey: second-best. Per-language and -task results for all models are in Appendix D.
@@ 5.1 Subnetwork Identification
Table 1 summarizes the number of language-specific neurons, corresponding trainable parameters, and the fraction of the model updated for each subnetwork within Llama-3.1-8B. Subnetwork sizes vary considerably: Welsh and Nepali subnetworks cover over 0.7% of model parameters, whereas Georgian and Macedonian subnetworks account for less than 0.3%. This variation might reflect differences in language complexity and the amount of language-specific signal captured by GlotCC. Subnetworks for Mistral-Nemo-12B and Aya-Expanse-8B (Appendix B) exhibit similar distributions and relative sizes.
Figure 2 and Appendix B.1 depict the distribution of the subnetwork neurons across layers for all 12 languages. We observe that most language-specific neurons are concentrated in the upper layers, and related languages exhibit moderate overlap in their neuron sets (such as Slovak and Slovenian, or Latvian and Lithuanian–see Appendix B.2 for details). These patterns align with prior findings for more high-resource languages from Gurgurov et al. (2025a). Overall, the weights associated with these subnetworks account for roughly 0.2% to 1% of the total model parameters, highlighting the efficiency of targeted adaptation.
@@ 5.2.1 Average Task Performance
Table 2 reports average downstream task performance across the 12 target languages for different fine-tuning configurations, along with results on general-purpose benchmarks for Llama-3.1-8B. Targeted sparse fine-tuning of language-specific subnetworks consistently yields the largest gains on target-language tasks, outperforming full-model, FFN-only, LoRA, IA3, and random-subset baselines, while fully preserving general capabilities as measured by MMLU and commonsense benchmarks (63.0 vs. 63.1 for the base model). For example, on FLORES, targeted subnetworks achieve average improvements of up to 5 BLEU points. This contrasts strongly with full-model and FFN-only fine-tuning, which not only fail to improve target-language performance but also substantially degrade general capabilities, a clear indicator of catastrophic forgetting and overfitting to the training distribution. The random neuron-subset and LoRA baselines show intermediate behavior: modest target-language gains but noticeable reductions in general performance, underscoring the benefit of principled subnetwork selection over arbitrary sparsity or auxiliary modules. IA3 does not degrade overall performance, but its improvements on target-language tasks are consistently small compared to targeted subnetwork fine-tuning.
5.2.2 Language-Wise Variation
As shown in Figure 3, the magnitude of improvement varies across languages. Languages such as Maltese and Estonian benefit the most from targeted fine-tuning, with average gains exceeding 8 points across all tasks, whereas Lithuanian and Slovak see only marginal improvements. A likely explanation is that higher-resource languages are already well represented in pre-training, limiting the marginal utility of additional fine-tuning data. In some cases, the model may even have encountered overlapping material during pre-training, reducing the benefit of adaptation (Lee et al., 2022b). By contrast, lower-resource languages, where the base model may have seen little to no data, gain substantially from subnetwork adaptation. This pattern is reflected in the moderate negative correlation between improvement scores and language resource availability (Appendix G), with improvements decreasing as resource levels increase.
The three models also exhibit distinct patterns in response to fine-tuning. Aya-Expanse-8B shows the most consistent improvements, with gains ranging from +2.1 (Slovak) to +21.5 (Maltese) and no negative results. Llama-3.1-8B achieves moderate gains across most languages (+1.7 to +8.8 points) but shows a small decrease for Georgian (-1.1). Mistral-Nemo-12B, despite its larger size, exhibits more variable behavior: strong gains for Maltese (+17.7) and Welsh (+10.4) are offset by declines for Latvian (-0.2), Nepali (-1.8), and Slovenian (-2.4). This pattern likely reflects differences in pre-training data composition and language capacity allocation. Notably, Aya-Expanse-8B–designed for multilingual capabilities–benefits most uniformly, suggesting models with stronger multilingual foundations are more amenable to subnetwork adaptation for low-resource languages.
The only observed decreases are for Georgian with Llama and Slovenian and Nepali with Mistral. Since both models still improve on these languages overall, and individual task results are mostly positive, these decreases may reflect that the allocated subnetwork is relatively small for effectively capturing language-specific knowledge for these languages, even if the absolute number of parameters is large. This suggests that some languages might benefit from different $\rho$% thresholds or alternative allocations across FFN components. Component-wise ablations (Appendix F) show that while fine-tuning all FFN projections performs best on average, some languages improve more by updating only specific components. We leave a systematic investigation of optimal subnetwork sizes and component-wise effects across languages to future work.
@@ 5.2.2 Language-Wise Variation
5.2.3 Training Dynamics
Figure 13(a) illustrates the training dynamics for Afrikaans. Sparse fine-tuning of target-language subnetworks generally converges faster and reaches lower validation loss than random-subset, FFN-only, full-model, LoRA, or IA3 fine-tuning. While full-model and FFN-only fine-tuning occasionally achieve lower loss on the target language (see Appendix C) due to the greater capacity of the updated networks, this is often accompanied by pronounced overfitting and a substantial drop in general capabilities.
5.2.4 Effect of Training Data Size
Figure 4 further examines the effect of varying training data size (50M, 100M, 150M, and 200M tokens) for a subset of seven languages (Macedonian, Latvian, Nepali, Slovak, Afrikaans, Maltese, and Lithuanian) using Llama-3.1-8B. For five out of seven languages, we observe that increasing data leads to improved performance, suggesting that additional data could further enhance results. This differs from the findings of Zhao et al. (2024), who fine-tune an ablation-identified subset of neurons using between 100 and 800 documents and report saturation after 400 documents. However, for Lithuanian and Maltese we similarly observe a decrease with larger training token sizes, which may indicate that the additional data is not helpful (Lee et al., 2022b) or the capacity of the identified subnetwork is insufficient to effectively utilize it. Due to computational constraints and the limited availability of data for some languages (e.g., Maltese and Afrikaans, with at most 100–150M tokens), we leave further exploration to future work.
Figure 4: Performance scores against various data sizes for a set of languages when sparsely fine-tuning Llama-3.1-8B.
6 Effects of Sparse Fine-Tuning
Beyond performance gains, we investigate the mechanisms underlying successful language adaptation. We analyze two aspects: weight changes induced by sparse fine-tuning and the evolution of cross-lingual representations.
(a) Up Projection
(b) Gate Projection
(c) Down Projection
@@ 6 Effects of Sparse Fine-Tuning
(a) Base
(b) Sparse Fine-tuning
@@
Figure 6: Cosine similarity for residual-stream activations of the original and sparsely fine-tuned Llama-3.1-8B between Maltese and reference languages (left, middle) and average across target languages (right). Full results in Appendix H.
@@ 6.1 Weight Change Analysis
Our analysis reveals a notable asymmetry between weight types within FFN layers. As shown in Figure 5, absolute weight changes for gate_proj and up_proj remain small and tightly clustered around zero across all layers, indicating highly precise adjustments. In contrast, down_proj exhibits substantially larger weight changes, particularly in later layers, where language-specific neurons are most concentrated.
[…]
6.2 Representation Drift Analysis
@@
To assess how targeted neuron fine-tuning affects cross-lingual representations, we measure whether parallel sentences become closer in embedding space after adaptation. Specifically, we compute cosine similarities between post-residual hidden states (mean-pooled across tokens) for Flores-200 sentences in the 12 target languages and 9 typologically diverse high-resource languages (Arabic, Chinese, English, French, German, Italian, Japanese, Russian, Spanish) across all layers (Figure 6). Results show consistent increases in cosine similarity after adaptation, particularly in later layers where early-layer representations already exhibit high similarity. As illustrated in Figure 6(c), targeted sparse fine-tuning achieves the strongest cross-lingual alignment (average cosine similarity of 90.5), outperforming all baselines including LoRA (87.5) and random-subset fine-tuning (87.7). Full fine-tuning, despite updating all parameters, achieves weaker alignment (87.3), likely due to representational drift from overfitting.
@@
These findings indicate that selectively updating identified neurons produces more language-agnostic representations in later layers, improving cross-lingual alignment without compromising model structure. This aligns with concurrent work showing improved alignment when manually manipulating language-relevant activations (Sundar et al., 2025), and suggests that sparse, targeted adaptation may be preferable to dense fine-tuning even when cross-lingual transfer is the primary goal.
7 Conclusion and Future Work
We present a framework for enhancing LLM capabilities in underrepresented languages via fine-tuning of language-specific subnetworks. Identifying language neurons and tuning their weights, our approach achieves substantial gains across 12 low- and mid-resource languages while preserving general-purpose performance. These results demonstrate that sparse, targeted adaptation can inject new language capabilities without catastrophic forgetting, addressing a key limitation of existing approaches. Mechanistic analyses reveal that adaptation concentrates in down-projection weights of later layers and improves cross-lingual alignment more effectively than dense fine-tuning. Our release of language neuron identifications for over 100 languages provides a practical resource for extending models to underrepresented languages. More broadly, our framework exemplifies actionable interpretability (Mosbach et al., 2024): we show that mechanistic insights can directly guide model enhancement, improving capabilities while preserving general performance. This highlights the importance of interpretability methods that enable optimization, not just post hoc analysis.
Looking forward, key directions include adaptively determining subnetwork size and composition based on language typology, data availability, and model scale, instead of fixed thresholds. Extending subnetwork adaptation to extremely low-resource languages and to other modalities such as vision and speech are also highly promising.
Impact Statement
This work advances the field of Machine Learning by enabling large language models to better handle underrepresented languages. Potential societal impacts include broader access to AI capabilities for speakers of low- and mid-resource languages and more equitable language technology, though risks such as misuse or biased outputs remain.